A car travels up a hill at a constant speed of 31 km/h and returns down the hill at a constant speed of 79 km/h. Calculate the average speed for the round trip.
Answer #1

Suppose uphill distance = x km

Time to uphill= distance/ speed = x/31 hours

Time to downhill = x/ 79 hours

Total time = x/31 + x/ 79 = 110x/ 2449

Total distance= 2x

So average speed= total distance/ total time

V =( 2x/110x)×2449 = 44.53 km/h